# A Network-Based Measure of Cosponsorship Influence on Bill Passing in the United States House of Representatives

### v1

@@ 4.2 Accounting for Political Parties
We account for the political parties of legislators by calculating the influence of legislators with Republicans and Democrats separately. We control for party affiliation by separately summing across the Democrat and Republican columns in $C_{\text{pass}}$ and $C_{\text{tot}}$. We then divide the Democrat and Republican sums by the number of Democrats and Republicans, respectively, in the House of Representatives at time $t$ to examine the relative influence that representative $i$ has with Democrats and Republicans. This yields tensors with entries
[…]
where $N_{\text{Dems}}$ and $N_{\text{{R}eps}}$, respectively, are the numbers of Democrats and Republicans in the House of Representatives at time $t$. The quantity $P[t,i]_{\text{{D}ems}}$ is the influence score of representative $i$ with Democrats at month $t$, and $P[t,i]_{\text{{R}eps}}$ is the influence score of representative $i$ with Republicans at month $t$.
@@ 4.3 Computing Influence at the Bill Level
To associate our measure of influence with bill-passage outcome, we examine influence at the bill level using time series of the mean and maximum cosponsor influence. The influence score of a bill aggregates the influence scores of its sponsors and cosponsors. We do this aggregation in two ways. In our first approach, we calculate a bill’s influence score as the mean of the influence scores of its cosponsors in the month that the bill was introduced. In our second approach, we calculate a bill’s influence score as the maximum of the influence scores of its cosponsors in the month that the bill is introduced. The two influence scores of bill $k$ are thus
@@
where $\mathcal{C}_{k}$ is the set of cosponsors of bill $k$, the number of cosponsors of bill $k$ is $|\mathcal{C}_{\text{k}}|$, and $t_{k}$ is the month that bill $k$ is introduced. These bill-level influence scores capture two different ideas. The mean assigns equal weights to each cosponsor, while ensuring that a large number of cosponsors does not artificially inflate a bill’s influence score. The maximum supposes that a bill’s most influential cosponsor carries it and thereby garners a large amount of support for it.
In Section 5, we compare our two bill-level influence scores and examine if either of them are indicative of whether a bill passes in the House of Representatives.
5 Influence as an Indicator of Success in Bill Passage
@@ 5.1 Visualizing Bill-Level Influence
We now illustrate how bill-level influence changes with time. In the left panels of Figure 2, we plot our bill-level influence scores (see Eq. (5)) as a function of time across nonoverlapping 4-month periods for bills that pass (orange crosses) and bills that fail (blue disks). We show the mean in Figure 2a and the maximum in Figure 2b. The corresponding histograms in the right panels show the distributions of the differences in the influence of passed and failed bills (i.e., the “relative-difference distributions” between passed and failed bills) in the House. We calculate this relative difference by determining the difference between the mean bill-level influence scores of passed and failed bills and dividing that difference by the mean influence score of failed bills. This yields the relative difference between passed and failed bills and thereby gives a normalized way to measure how much the influence of passed bills exceeds the influence of failed bills.
@@
We use 4-month time periods as our (nonoverlapping) aggregation time windows for two key reasons. First, there are some months (particularly towards the end of each year) in which the House of Representatives passes very few bills. Aggregating data over a few months thus smooths out temporally localized fluctuations. Second, in the employed data, bills took a median time of 98 days (i.e., 3.2 months) and a mean time of 148 days (i.e., 4.9 months) between the time that they were introduced to the time that they passed in the House. Bills took a longer time to pass in the Senate, with a median time of 148 days (i.e., 4.9 months) and a mean time of 196 days (i.e., 6.5 months). Our choice of 4 months as an aggregation time is informed by the median and mean passage times in the House. Naturally, other choices are also viable, and future work can explore the effect of aggregating influence over different amounts of time. One can also consider overlapping time windows.
@@
Figure 2: Bill-influence time series and the corresponding distributions of relative differences between influence scores of bills that pass and fail in the House of Representatives. In (a), we show bill-level influence that we calculate using the mean of a bill’s cosponsors’ influence scores. In (b), we show bill-level influence that we calculate using the maximum of a bill’s cosponsors’ influence scores. In the left panels, we show bill-level influence for 10 years, which we aggregate into nonoverlapping 4-month periods. Each point is the mean influence score of bills in the associated 4-month period starting on the date on the horizontal axis. The orange crosses are bills that fail in the House, and the blue disks are bills that pass in the House. The error bar on each marker indicates the standard error. Our computation of influence incorporates an exponential decay with a half-life of 6 months. In the right panels, we plot the distributions of the relative differences between the mean influence scores of passed and failed bills. The horizontal axis gives the relative differences between the mean influence scores of bills that pass and fail in the House of Representatives. The vertical axis gives the counts of these relative differences.
@@
As we can see in Figure 2, our notions of influence separate passed bills and failed bills, with passed bills having consistently larger influence scores than failed bills for all 4-month periods between 2009 and 2019 (i.e., Congresses 111–115). Therefore, our notion of influence is a relevant indicator of bill passage. In Figure 2, we also observe that calculating a bill-level influence score as the maximum of the cosponsors’ influences better differentiates between passed and failed bills than calculating bill-level influence as the mean of cosponsors’ influences. See the relative-difference distributions in the right panels. When we use a half-life of 6 months, the mean of the relative-difference distribution between bill-level influence scores of passed and failed bills is $0.248$ for maximum cosponsor influence and $0.139$ for mean cosponsor influence. We obtain qualitatively similar results for half-lives of 12 months and 24 months. For a half-life of 12 months, the relative difference between influence scores is $0.223$ for the maximum influence score and $0.122$ for the mean influence score; for a half-life of 24 months, the relative difference between influence scores is $0.203$ for the maximum and $0.109$ for the mean. See Section 5.2 for more details. Accordingly, we conclude that, on average, a bill is more likely to pass in the House if a single very influential representative cosponsors it than if several moderately influential representatives all cosponsor it. This idea relates to opinion leadership and the ‘influentials hypothesis’, which suggests that a few people tend to be the predominant influencers of the majority of a population and the formation of public opinion [48]. This phenomenon and related (but somewhat different) phenomena are relevant in various decision-making situations, such as voting and purchasing [49, 50].
@@
There are many other possible ways — beyond using simply the maximum or mean of cosponsor influence — to compute a bill-level influence score from a legislator-level influence score. One possibility is to combine contributions of very influential legislators with several moderately influential legislators by employing thresholded (by, e.g., averaging the influences of all sufficiently influential legislators) or weighted means of cosponsor influence.
@@ 5.2 Effect of Influence Time Scale
Our results in Section 5 suggest that using cosponsorship information to construct an influence measure of bills is informative for studying whether or not a given bill is likely to pass. In this subsection, we examine the effect of the influence decay half-life on how indicative our bill-level influence score (with maximum cosponsorship influence) is on bill passage.
@@
Figure 3: Distribution of the relative differences in influence between passed and failed bills for half-lives of 6, 12, and 24 months. We calculate the influence using the maximum influence score of a bill’s cosponsors, and we calculate the relative difference by normalizing the difference between influence scores of the passed and failed bills by the influence scores of the failed bills.
In Figure 3, we show the effect of the half-life of a legislator’s influence decay on our ability to differentiate between passed and failed bills. For all three half-lives, we are similarly successful at distinguishing between passed and failed bills. The relative differences between the influence scores are 0.248, 0.223, and 0.203 for half-lives of 6, 12, and 24 months, respectively. The standard errors are 0.0139, 0.0119, and 0.0107, respectively. The mean relative-differences for the half-lives differ from each other by more than one standard error, indicating that half-life is a meaningful parameter. For slower decay rates, we obtain smaller differences between the influence scores of passed and failed bills.
@@ 5.3 Comparison of the Results for our Influence Measure to those for Eigenvector Centrality
It is common to use centrality measures to examine the importances of the nodes of a network [18, 51]. Therefore, in principle, one can calculate centralities to attempt to measure the influence of legislators. One can also interpret our legislator influence scores as centrality measures.
We now calculate the eigenvector centrality of each node (i.e., legislator), which measures a node’s importance based on the importance of its neighboring nodes in a network, and find that it is not a strong indicator of bill passage. In Appendix A, we consider closeness centrality and strength centrality (i.e., weighted degree centrality), which are also less indicative than our influence measures of whether or not a bill passes in the House.
The eigenvector centrality of a node is given by its associated entry of the leading eigenvector of the adjacency matrix of a graph [18]. The eigenvector centrality of a node of a network is large when its adjacent nodes have large eigenvector centralities. Therefore, the eigenvector centrality of a representative can be large even when their degree is not large. To calculate eigenvector centrality, we first divide the tensor entry $C_{\text{pass}}[t,i,j]$ (1) (i.e., the decayed number at time $t$ of passed bills that include representatives $i$ and $j$ as cosponsors) by the tensor entry $C_{\text{tot}}[t,i,j]$ (1) (i.e., the decayed number at time $t$ of introduced bills that include representatives $i$ and $j$ as cosponsors) to yield a tensor $D$ with entries
@@
We then determine the largest connected component $D[t]_{\text{LCC}}$ of the network that corresponds to the month-$t$ adjacency matrix of the tensor $D$. We obtain the eigenvector centrality of representative $i$ at month $t$ by solving
@@
Figure 4: Eigenvector-centrality time series and the corresponding distributions of relative differences between the eigenvector centralities of bills that pass and fail in the House of Representatives. In (a), we show the bill-level eigenvector centrality that we calculate using the mean of the eigenvector centralities of a bill’s cosponsors. In (b), we show the bill-level eigenvector centrality that we calculate using the maximum of the eigenvector centralities of a bill’s cosponsors. In the left panels, we show bill-level influence for 10 years, which we aggregate into nonoverlapping 4-month periods. Each point is the mean eigenvector centrality for bills in the associated 4-month period starting on the date on the horizontal axis. The orange crosses are bills that fail in the House, and the blue disks are bills that pass in the House. The error bar on each marker indicates the standard error. Our computation of eigenvector centrality incorporates an exponential decay with a half-life of 6 months. In the right panels, we plot the distributions of the relative differences between the eigenvector centralities of passed and failed bills. The horizontal axis gives the relative differences between the mean eigenvector centralities of bills that pass and fail in the House of Representatives. The vertical axis gives the counts of these relative differences.
By comparing Figure 4 to Figure 2 (see Section 5.1), we see that eigenvector centrality is a much worse indicator of bill-passage success than our influence measure. The mean relative difference between passed and failed bills for the mean eigenvector centrality is $0.004$, and the mean relative different for the maximum eigenvector centrality is $0.002$. These numbers are much smaller than the mean relative differences for mean influence ($0.139$) and maximum influence ($0.248$).
@@ 6 Conclusions and Discussion
We studied the effect of bill cosponsorship relationships between legislators on bill-passage success in the US House of Representatives. We introduced a notion of influence in which legislators gain influence with each other whenever a bill that they cosponsor passes. This legislator influence, which we normalize by how often legislators cosponsor a bill, accumulates for each cosponsored bill that passes and decays exponentially with time. We then examined how much our influence notion correlates with the passage rate of subsequent bills.
@@
Our calculations suggest that our influence measure, which we obtain by analyzing bill cosponsorship relationships through time, is informative about the voting outcomes of bills. For example, based on our influence measure, cosponsorship by a single very influential representative appears to be a better indicator of bill-passage success in the US House of Representatives than cosponsorship by several moderately influential representatives. However, our work has important limitations. In particular, our notion of legislator influence is simplistic, and obviously it does not perfectly capture the social nuances of influence. For instance, we treat all legislators identically, but clearly there are many differences between different legislators. For example, they are subject to different external influences and may have rather different goals. Bill cosponsorship information does seem to yield a proxy of legislator influence, but it is only a proxy.
@@
There are a variety of extensions to pursue in future work. For example, it is desirable to account for factors (such as bill topic and committee assignment) beyond legislator influence that affect bill outcome. This will help isolate the effects of legislator influence on bill-passage success. There are also other data sources beyond bill cosponsorship that encapsulate aspects of the relationships between legislators. It is also worth investigating how much our simplistic notion of influence correlates with other features of legislators, such as personal demographics, seniority, and funding from interest groups. It also seems interesting to examine polyadic influence, as perhaps multiple legislators together influence other legislators.

### v2

@@ 4.2 Accounting for Political Parties
We account for legislators’ political parties by calculating the influence of legislators with Republicans and Democrats separately. We control for party affiliation by separately summing across the Democrat and Republican columns in $C_{\text{pass}}$ and $C_{\text{tot}}$. We then divide the Democrat and Republican sums by the number of Democrats and number of Republicans, respectively, in the House of Representatives at time $t$ to examine the relative influence that representative $i$ has with Democrats and Republicans. This yields tensors with entries
$P[t,i]_{\text{{D}ems}}$ $=\frac{1}{N_{\text{Dems}}}\cdot\frac{\sum_{d}C[t,i,d]_{\text{pass}}}{\sum_{d}C[t,i,d]_{\text{tot}}}\,,\,\,\,d\in{\text{\{Democrats\}}}\,,\,\,{N}_{\text{Dems}}\,=|{\text{\{Democrats\}}}|\,$,
$\vskip 17.07164ptP[t,i]_{\text{{R}eps}}$ $=\frac{1}{N_{\text{Reps}}}\cdot\frac{\sum_{d}C[t,i,r]_{\text{pass}}}{\sum_{d}C[t,i,r]_{\text{tot}}}\,,\,\,\,r\in{\text{\{Republicans\}}}\,,\,\,{N}_{\text{Reps}}\,=|{\text{\{Republicans\}}}|\,$, (3)
where $N_{\text{Dems}}$ and $N_{\text{{R}eps}}$, respectively, are the numbers of Democrats and Republicans in the House of Representatives at time $t$. The quantity $P[t,i]_{\text{{D}ems}}$ is the influence score of representative $i$ with Democrats in month $t$, and $P[t,i]_{\text{{R}eps}}$ is the influence score of representative $i$ with Republicans in month $t$.
@@
Figure 2: Distribution of individual legislators in the US House of Representative in the final month of each Congress.
As we can see in Figure 2, the distribution of the influence scores of individual legislators in the House of Representative ranges between 0 and $0.008$. Perhaps unintutively, the mean of this distribution does not increase with time. In fact, the distribution mean seems to be smallest in Congresses 114 and 115. We hypothesis that this decrease in the distribution mean from the mean in Congresses 111–113 may be a signal of less collaboration between representatives in Congresses 114–115 than in the prior three Congresses. Congresses 114–115 coincide with the years 2015–2019, which saw a documented increase in political polarization in the US Congress [48].
@@ 4.3 Computing Influence at the Bill Level
To associate our measure of influence with bill-passage outcome, we examine influence at the bill level using time series of the mean and maximum cosponsor influence. The influence score of a bill aggregates the influence scores of its sponsors and cosponsors. We do this aggregation in two ways. In our first approach, we calculate a bill’s influence score as the mean of the influence scores of its cosponsors in the month that the bill is introduced. In our second approach, we calculate a bill’s influence score as the maximum of the influence scores of its cosponsors in the month that the bill is introduced. The two influence scores of bill $k$ are thus
@@
where $\mathcal{C}_{k}$ is the set of cosponsors of bill $k$, the number of cosponsors of bill $k$ is $|\mathcal{C}_{\text{k}}|$, and $t_{k}$ is the month that bill $k$ is introduced.
Our two bill-level influence scores capture two different ideas of legislator influence. The mean assigns equal weights to each cosponsor, while ensuring that a large number of cosponsors does not artificially inflate a bill’s influence score. The maximum supposes that a bill’s most influential cosponsor carries it and thereby garners a large amount of support for it.
Because we determine bill influence from legislator influence, which depends on the bill-passage outcome of many bills, whether or not a specific bill passes exerts only a small influence on its own influence score. We aggregate our results at a monthly level, whereas legislators vote on a bill on a specific day. In our aggregation, influence must either (1) include a bill in its own calculation or (2) not consider influence that arises from an earlier day of the same month. The first option can only change a numerator in Eq. (5) by a maximum of 1 bill and the second option can yield a much more substantial change, so we use the first option. Additionally, these numerators are fairly large, so an increase of it by 1 bill has little effect on the calculation.
5 Bill-Level Influence as an Indicator of Success in Bill Passage
We now compare our two bill-level influence scores and examine if either of them is indicative of whether a bill passes in the House of Representatives.
@@ 5.1 Visualizing Bill-Level Influence
We now illustrate how bill-level influence changes with time. These changes in bill-level influence reflect aggregate changes in cosponsor influence with time (see Eq. (5)).
@@
In the left panels of Figure 3, we plot our bill-level influence scores as a function of time across nonoverlapping 4-month periods for bills that pass (orange crosses) and bills that fail (blue disks). We show bill-level influence scores from mean cosponsor influence in Figure 3a, and we show the bill-level influence scores from maximum cosponsor influence in Figure 3b. The corresponding histograms in the right panels show the distributions of the differences in mean bill-level influence between passed and failed bills (i.e., the ‘relative-difference distributions’ between passed and failed bills) in the House. For both mean cosponsor influence and maximum cosponsoe influence, we calculate these relative differences by determining the difference between the mean bill-level influence scores of passed and failed bills and dividing that difference by the mean influence score of failed bills. This yields relative differences between passed and failed bills, and it thereby gives a normalized way to measure how much the bill-level influence of passed bills exceeds the bill-level influence of failed bills.
@@
We use 4-month time periods as our (nonoverlapping) aggregation time windows for two key reasons. First, there are some months (especially near the end of each year) in which the House of Representatives passes very few bills. Aggregating data over a few months thus smooths out temporally localized fluctuations. Second, in the employed data, bills took a median time of 98 days (i.e., 3.2 months) and a mean time of 148 days (i.e., 4.9 months) between the time that they were introduced to the time that they passed in the House. Bills took a longer time to pass in the Senate, with a median time of 148 days (i.e., 4.9 months) and a mean time of 196 days (i.e., 6.5 months). Our choice of 4 months as an aggregation time is informed by the median and mean passage times in the House. Naturally, other choices are also viable, and future work can explore the effect of aggregating influence over different amounts of time. One can also consider overlapping time windows.
@@
Figure 3: Bill-influence time series and the corresponding distributions of relative differences between influence scores of bills that pass and fail in the House of Representatives. In (a), we show bill-level influence that we calculate using the mean of a bill’s cosponsors’ influence scores. In (b), we show bill-level influence that we calculate using the maximum of a bill’s cosponsors’ influence scores. In the left panels, we show bill-level influence for 10 years, which we aggregate into nonoverlapping 4-month periods. Each point is the mean influence score of bills in the associated 4-month period starting on the date on the horizontal axis. The values of the bill-level influence scores range between 0 and $1$. The orange crosses are bills that fail in the House, and the blue disks are bills that pass in the House. The error bar on each marker indicates the standard error. Our computation of influence scores incorporates an exponential decay with a half-life of 6 months. In the right panels, we plot the distributions of the relative differences between the mean influence scores of passed and failed bills. The horizontal axis gives the relative differences between the mean influence scores of bills that pass and fail in the House of Representatives. The vertical axis gives the counts of these relative differences.
@@
In Figure 3, we see that our notions of bill-level influence separate passed bills and failed bills, with passed bills having consistently larger influence scores than failed bills for all 4-month periods between 2009 and 2019 (i.e., Congresses 111–115). Therefore, our bill-level notion of influence is a relevant indicator of bill passage. In Figure 3, we also see that calculating a bill-level influence score as the maximum of cosponsors’ influences better differentiates between passed and failed bills than calculating bill-level influence as the mean of cosponsors’ influences. In particular, observe the relative-difference distributions in the right panels. When we use a half-life of 6 months, the mean of the relative-difference distribution between bill-level influence scores of passed and failed bills is $0.248$ for maximum cosponsor influence and $0.139$ for mean cosponsor influence. We obtain qualitatively similar results for half-lives of 12 months and 24 months. For a half-life of 12 months, the relative difference between influence scores is $0.223$ for the maximum influence score and $0.122$ for the mean influence score. For a half-life of 24 months, the relative difference between influence scores is $0.203$ for the maximum and $0.109$ for the mean. See Section 5.2 for more details. Accordingly, we conclude that, on average, a bill is more likely to pass in the House if a single very influential representative cosponsors it than if several moderately influential representatives all cosponsor it. This idea relates to opinion leadership and the ‘influentials hypothesis’, which suggests that a few people tend to be the predominant influencers of the majority of a population and the formation of public opinion [49]. This phenomenon and related (but somewhat different) phenomena are relevant in various decision-making situations, such as voting and purchasing [50, 51]. We also observe a steep decline in both the mean and maximum bill-level influence scores with time, most noticeably starting in about 2015. We hypothesize that this decline is correlated with the increase in political polarization in the US during this time [48].
@@
There are many other possible ways — beyond using simply the mean or maximum of the cosponsor influence — to compute a bill-level influence score from a legislator-level influence score. One possibility is to combine contributions of very influential legislators with several moderately influential legislators by employing thresholded means (e.g., by averaging the influence of all sufficiently influential legislators) or weighted means of cosponsor influence.
@@ 5.2 Effect of Influence Time Scale
Our results in Section 5.1 suggest that using cosponsorship information to construct an influence measure of bills is informative for studying whether or not a given bill is likely to pass. In this subsection, we examine the effect of the half-life of the influence decay on how indicative our bill-level influence score (with maximum cosponsorship influence) is on bill passage.
@@
Figure 4: Distribution of the relative differences in influence between passed and failed bills for half-lives of 6, 12, and 24 months. We calculate the influence using the maximum influence score of a bill’s cosponsors, and we calculate the relative difference by normalizing the difference between the influence scores of passed and failed bills by the influence scores of the failed bills.
In Figure 4, we show the effect of the half-life of a legislator’s influence decay on our ability to differentiate between passed and failed bills. For all three half-lives, we are similarly successful at distinguishing between passed and failed bills. The relative differences between the influence scores are 0.248, 0.223, and 0.203 for half-lives of 6, 12, and 24 months, respectively. The standard errors are 0.0139, 0.0119, and 0.0107, respectively. For slower decay rates, we obtain smaller differences between the influence scores of passed and failed bills. However, from a practical point of view, the differences in these influence scores are only slightly different for different half-lives. Therefore, choosing any half-life between 6 months and 24 months gives a similar ability to distinguish between passed and failed bills.
@@ 5.3 Comparison of the Results for our Influence Measure to those for Eigenvector Centrality
It is common to use centrality measures to examine the importances of the nodes of a network [52, 18, 53]. Therefore, in principle, one can calculate centralities to attempt to measure the influence of legislators. One can also interpret our legislator influence scores as centrality measures.
We now calculate the eigenvector centrality of each node (i.e., legislator). Eigenvector centrality quantifies the importance of a node based on the importances of its neighboring nodes in a network. Based on our calculations, we find that it is not a strong indicator of bill passage. In Appendix A, we consider closeness centrality and strength centrality (i.e., weighted degree centrality), which are also less indicative than our influence measure of whether or not a bill passes in the House.
The eigenvector centrality of a node is given by its associated entry of the leading eigenvector of the adjacency matrix of a graph [18]. The eigenvector centrality of a node of a network is large when its adjacent nodes have large eigenvector centralities. Therefore, the eigenvector centrality of a node can be large even when its degree is not large. To calculate eigenvector centrality, we first divide the tensor entry $C_{\text{pass}}[t,i,j]$ (1) (i.e., the decayed number in month $t$ of passed bills that include representatives $i$ and $j$ as cosponsors) by the tensor entry $C_{\text{tot}}[t,i,j]$ (1) (i.e., the decayed number in month $t$ of introduced bills that include representatives $i$ and $j$ as cosponsors) to yield a tensor $D$ with entries
@@
We then determine the largest connected component $D[t]_{\text{LCC}}$ of the network that corresponds to the month-$t$ adjacency matrix of the tensor $D$. We obtain the eigenvector centrality of representative $i$ in month $t$ by solving
@@
Figure 5: Eigenvector-centrality time series and the corresponding distributions of relative differences between the eigenvector centralities of bills that pass and fail in the House of Representatives. In (a), we show the bill-level eigenvector centralities that we calculate using the mean of the eigenvector centralities of a bill’s cosponsors. In (b), we show the bill-level eigenvector centralities that we calculate using the maximum of the eigenvector centralities of a bill’s cosponsors. In the left panels, we show bill-level eigenvector centralities for 10 years, which we aggregate into nonoverlapping 4-month periods. Each point is the mean eigenvector centrality for bills in the associated 4-month period starting on the date on the horizontal axis. The values of the eigenvector centralities range between 0 and $1$. The orange crosses are bills that fail in the House, and the blue disks are bills that pass in the House. The error bar on each marker indicates the standard error. Our computation of eigenvector centralities incorporates an exponential decay with a half-life of 6 months. In the right panels, we plot the distributions of the relative differences between the eigenvector centralities of passed and failed bills. The horizontal axis gives the relative differences between the mean eigenvector centralities of bills that pass and fail in the House of Representatives. The vertical axis gives the counts of these relative differences.
By comparing Figure 5 to Figure 3 (see Section 5.1), we see that eigenvector centrality is a much worse indicator of bill-passage success than our influence measure. The mean relative difference between passed and failed bills for the mean eigenvector centrality is $0.004$, and the mean relative difference for the maximum eigenvector centrality is $0.002$. These numbers are much smaller than the mean relative differences for mean influence ($0.139$) and maximum influence ($0.248$).
@@ 6 Conclusions and Discussion
We studied the effect of bill cosponsorship relationships between legislators on bill-passage success in the US House of Representatives. To do this, we introduced a notion of influence in which legislators gain influence with each other whenever a bill that they cosponsor passes. This legislator influence, which we normalize by how often legislators cosponsor a bill, accumulates for each cosponsored bill that passes and decays exponentially with time. We leveraged our notion of legislator influence to quantify bill-level influence and study correlations between influence and the bill-passage rate.
@@
Our calculations suggest that our influence notion, which we obtain by analyzing bill cosponsorship relationships through time, is informative about the voting outcomes of bills. For example, based on our influence measure, cosponsorship by a single very influential representative appears to be a better indicator of bill-passage success in the US House of Representatives than cosponsorship by several moderately influential representatives. We thus conclude that bill cosponsorship information yields a proxy of legislator influence, but it is only a proxy.
Our work has important limitations. In particular, our notion of legislator influence is simplistic, and obviously it does not perfectly capture the social nuances of influence. For instance, we treated all legislators identically, but clearly there are many differences between different legislators. Different legislators are subject to different external influences and may have rather different goals. Another notable limitation of our analysis is that we focused on aggregated results. We purposely refained from commenting on legislator-level observations (such as which legislators are most and least influential), as we have much greater confidence in aggregate results than in individual legislator-level results. It seems valuable to build on our investigation by probing the influence dynamics of individual legislators and notable groups of legislators [54].
@@
There are many extensions of our study to pursue in future work. For example, it is desirable to account for factors (such as bill topics and committee assignments) beyond legislator influence that affect voting outcomes. Examining the impact of such factors will help isolate the effects of legislator influence on bill-passage success. It is also worth investigating how much our simplistic notion of influence correlates with other features of legislators, such as personal demographics, seniority, and funding from interest groups. Additionally, it is worthwhile to examine what major political and/or social events may have influenced the temporal trends that we observed in our influence measure.
It is also important to study other links between legislator networks and their voting and cosponsorship choices. Other measures encapsulate important aspects of the relationships between legislators. One notable measure from the political-science literature is the legislative-effectiveness score (LES) [55]. Like our influence notion, LES considers bill-voting outcomes and cosponsorship behavior. However, it differs from our approach by looking at more granular stages of the legislative process, characterizing bills by their importance, and not including a decay factor. To conduct a deeper investigation of influence dynamics in the US Congress, it is relevant to compare our legislator influence scores with LESs. It is also worthwhile to examine some of the numerous measures from network science [52]. In our analysis, we compared the results from our influence notion to results from three conventional centrality measures (specifically, eigenvector, closeness, and strength centralities), and we observed that those traditional centralities do not correlate strongly with bill-passage success. There are a suite of other centralities and additional network measures, and naturally one can examine other measures — such as network efficiency [56], which is concerned with how well information propagates in a network — to see how much they correlate with bill-passage success. It also seems interesting to examine polyadic influence (e.g., see [57]), as perhaps multiple legislators together (e.g., legislator groups like the Squad [58]) influence other legislators.